OpenStudy (anonymous):

esitmate :8 1/8 * 3 5/6= exact:8 1/8 * 3 5/6=

OpenStudy (anonymous):

8 1/8 * 3 5/6 = 15/6

OpenStudy (anonymous):

estimate or exact

Parth (parthkohli):

Estimate: 8 * 4 = 32 Exact = 8 1/8 * 2 5/6 = 15/6

OpenStudy (anonymous):

exact more simplificat: 5/2

OpenStudy (anonymous):

15/6 = 5/2

Parth (parthkohli):

Oops. Estimate is 8 * 3 = 24

OpenStudy (anonymous):

final-> 15/6 exact-> 5/2 decimal->2,5